# When Does Learning Renormalize? Sufficient Conditions for Power–Law Spectral Dynamics

### v3

@@ 1 Introduction
Empirical power–law scaling has emerged as one of the most robust and reproducible phenomena in modern deep learning. Across model families, tasks, and training regimes, performance metrics such as loss, error, or perplexity often exhibit smooth power–law dependence on model size, dataset size, or compute budget (hestness2017deep; kaplan2020scaling; henighan2020scaling; hoffmann2022training). More recently, scaling behavior has been observed to persist beyond classical regimes, extending to precision, data pruning, and architectural interventions (sorscher2022beyond; kumar2024scaling). Despite its empirical ubiquity, the theoretical origins of these power laws remain only partially understood.
A growing body of work has sought to explain neural scaling laws through kernel limits and mean–field approximations (jacot2018neural; lee2019wide; yang2021tensor), as well as through dynamical models that interpolate between lazy and feature–learning regimes (bordelon2024dynamical; bordelon2023feature; domine2024lazy). While these approaches provide valuable insight into specific limits or architectures, they do not by themselves explain why power–law behavior arises so broadly, nor under what conditions it should be expected to fail. In particular, empirical evidence clearly indicates that scaling laws do not hold universally: depending on architecture, optimization stability, data distribution, or training regime, scaling behavior may degrade, break, or transition between distinct regimes (sorscher2022beyond; keskar2017large; ghorbani2019investigation).
In recent work, the Generalized Resolution–Shell Dynamics (GRSD) framework was proposed as a spectral and operator–theoretic description of learning dynamics in deep networks (zhang2025generalized). GRSD formulates learning as an energy transport process across logarithmic spectral shells, leading to a coarse–grained, one–dimensional conservation law for error energy. Within this framework, empirical neural scaling laws correspond to a particularly simple form of the renormalized shell velocity: a power–law function of the spectral coordinate. However, in GRSD this power–law form is not automatic. Rather, it reflects a nontrivial structural property of the learning process—namely, that the shell dynamics admits a renormalizable closure in the sense familiar from statistical physics and turbulence theory (kadanoff1966scaling; wilson1983renormalization; kolmogorov1995turbulence; forster1977large; spohn2012large).
@@ 2 Generalized Resolution–Shell Dynamics
In this section we briefly review the Generalized Resolution–Shell Dynamics (GRSD) framework, which provides the dynamical and spectral setting for the results of this paper. Our presentation is intentionally concise and focuses only on the aspects of GRSD that are required to formulate and analyze power–law renormalizability. We refer to (zhang2025operator) for a detailed and comprehensive treatment.
@@ 2.1 Spectral shells and shell energies
which plays the role of a time-dependent kernel governing error dynamics. Operators of this form have been extensively studied in learning theory and operator-based analyses of neural networks (rosasco2010learning; koltchinskii2017normal).
@@ 2.2 Renormalized shell dynamics
where $J(\lambda,t)$ is a spectral flux density describing energy transport across scales, and $D(\lambda,t)$ collects dissipative contributions. This equation is the learning-theoretic analogue of energy cascade equations in turbulence and large-scale interacting systems (kolmogorov1995turbulence; forster1977large; spohn2012large).
@@ 3.1 Learning configurations and Jacobian dynamics
which governs the instantaneous learning dynamics in function space. Operator evolutions of this form arise naturally in both kernel limits and feature-learning regimes (jacot2018neural; lee2019wide; bordelon2024dynamical; zhang2025operator).
@@ 3.2 Statement of sufficient conditions
Assumptions of this type—uniformly bounded Jacobians/gradients along the optimization trajectory, sometimes supplemented by smoothness controls—are standard in analyses of overparameterized learning and trainability; see, e.g., bounded-Jacobian conditions in terjek2022framework and Jacobian-spectrum stability via dynamical isometry in pennington2017resurrecting; tarnowski2019dynamical.
@@ 3.3 Main theorem
Fix a learning configuration and consider the induced GRSD shell dynamics defined on logarithmic spectral shells. Suppose Conditions 14 hold on a training horizon $t\in[t_{0},T]$. Assume in addition the standard GRSD structural properties: (i) antisymmetry of intra–shell couplings, so that shell–internal transfers cancel in the energy balance; and (ii) renormalizable closure, whereby inter–shell dynamics can be expressed up to a controlled remainder in terms of boundary fluxes between neighboring shells.
@@
then the renormalized GRSD velocity field is uniquely constrained to
$v(\lambda,t)=\frac{c_{0}}{t}\lambda,t\in[t_{0},T]$
@@
When a learning–rate schedule $\eta(t)$ is present, the parameter dynamics take the form $\dot{\theta}=-\eta(t)\nabla_{\theta}\mathcal{L}$ and can be mapped to the above gradient–flow form by introducing a reparameterized time lifunctional
@@ 3.4 Proof overview
We briefly outline the structure of the proof; technical details are deferred to the appendix. First, Conditions 13 imply that locality and summable incoherence of Jacobian blocks propagate along the training trajectory. This yields a one-dimensional shell conservation law with boundary fluxes governing inter–shell energy transport. Second, Condition 4 implies homogeneity of shell-averaged energy and flux densities under spectral rescaling. Finally, homogeneity of the ratio $v(\lambda,t)=J(\lambda,t)/\varepsilon(\lambda,t)$ enforces a power–law functional form. Detailed proof is included in Appendix A.2
@@ 4.2 Residual learning implies log–shift invariance
The residual form $A_{k}=I+\varepsilon G_{\ell}$ induces additive log–spectral increments across depth. Under the absence of depth–wise parameter sharing and the controlled path regularity of Condition 3, these increments define a stationary Markov–additive process on spectral directions. Two mechanisms govern the depth threshold. First, Berry–Esseen type bounds bolthausen1982berryesseen yield self–averaging of accumulated log–spectral increments once $L\gtrsim\varepsilon^{-2}\eta^{-2}$. Second, the near–identity residual perturbations induce quantitative mixing on the projective space, leading to isotropization of spectral directions on a scale $L\gtrsim\varepsilon^{-2}\log(1/\eta)$. Beyond the larger of these two thresholds, bin–averaged coupling statistics become translation–invariant in logarithmic scale. A complete proof is provided in Appendix A.3. ∎
Theorem 2 identifies residual learning as a mechanism by which individual layerwise Jacobian contributions become asymptotically free of absolute scale information once depth exceeds a finite threshold. Its proof relies in an essential way on the near–identity residual parameterization and depth–wise statistical stationarity. The emergence of directionally homogeneous behavior under repeated small random perturbations is consistent with classical results on random matrix products and Markovian mixing on projective spaces (furstenberg1960products; levin2009markov). As a result, the argument does not directly extend to architectures with explicit gating or depth–dependent weighting, nor to general non–skip architectures. This limitation should not be interpreted as ruling out scaling behavior in such models; rather, it indicates that different analytical tools would be required to verify Condition 4 outside the residual setting.
@@ 5.1 Boundedness of the gradient computation graph
A canonical example of a bounded computation graph is a directed acyclic graph (DAG) abadi2016tensorflow; paszke2019pytorch, such as those induced by standard feedforward networks, residual architectures, or transformers. In these settings, gradient propagation is naturally confined by the depth of the graph, and the influence of any parameter update cannot spread arbitrarily far in an infinitesimal time step. As a result, the induced Jacobian evolution is uniformly bounded in the sense required by Condition 1.
By contrast, recurrent architectures with unconstrained feedback loops may violate this boundedness condition. In particular, when cyclic structures contribute divergent amplification—such as repeated multiplication by unstable recurrent weights—the Jacobian evolution may accumulate unbounded long-range couplings. Classical recurrent neural networks (RNN) hopfield1982neural and certain long short-term memory (LSTM) hochreiter1997long configurations fall into this category when their recurrent dynamics is not properly controlled. In such cases, the GRSD coarse-graining procedure may fail, as infinitesimal updates can induce global spectral reorganization.
Importantly, the presence of cycles alone does not preclude boundedness. Recurrent or iterative architectures may still satisfy Condition 1 provided that the contribution of cycles is contractive or effectively damped. For example, when recurrent interactions decay exponentially gu2021efficiently; smith2022simplified; gu2024mamba; peng2023rwkv; xu2025bringing, the cumulative influence of long feedback paths remains finite, and Jacobian evolution can be bounded despite the existence of loops. From the GRSD perspective, such architectures behave similarly to bounded-depth computation graphs at the level of infinitesimal learning dynamics. We provide a precise proposition and a proof sketch for this case in Appendix 2
@@ 5.2 Initial functional incoherence
Condition 2 constrains the geometry of the Jacobian at initialization. It requires that correlations between distant Jacobian blocks, measured through operator inner products, decay sufficiently fast so that their tail is summable. Modern zero-mean random i.i.d. initializations glorot2010understanding; he2015delving; klambauer2017self commonly used in deep learning naturally satisfy the condition of initial gradient incoherence, in the sense that gradient components associated with distinct functional modes are uncorrelated at initialization.
@@ 5.3 Controlled Jacobian evolution
This assumption should be viewed as an abstraction of common stability-inducing practices in modern deep learning, such as normalization, residual connections, and conservative optimization schedules (terjek2022framework; pennington2017resurrecting; tarnowski2019dynamical). However, the condition itself is agnostic to the specific optimization algorithm and does not rely on stochasticity or noise-induced regularization.
@@ 5.4 Log–shift invariance and rigidity of renormalized dynamics
Within the GRSD framework, log–shift invariance alone enforces homogeneity of the renormalized shell equations, but does not by itself fix the functional form of the shell velocity. The emergence of a power–law velocity is instead a rigidity phenomenon: when log–shift invariance is combined with the intrinsic time–rescaling covariance of gradient flow, all admissible functional forms except pure power laws are excluded. Such rigidity phenomena, in which scale invariance combined with consistency constraints restrict admissible large-scale behavior to power laws, are well known in renormalization-group analyses of critical phenomena and turbulent transport (kadanoff1966scaling; wilson1983renormalization; forster1977large). If Condition 4 fails, the shell dynamics may exhibit scale–dependent drift, moving spectral reference points, or broken scaling behavior, even if the remaining conditions are satisfied.
@@ 6.1 Multilayer perceptrons
Multilayer perceptrons (MLPs) rosenblatt1957perceptron provide the simplest setting in which the sufficient conditions can be examined. Their feedforward structure naturally induces locality in the gradient computation graph, consistent with Condition 1. Moreover, standard random initialization schemes typically lead to weak functional correlations between distant layers in sufficiently wide networks, aligning with Condition 2 in early training regimes.
@@ 6.2 Convolutional networks
An important exception arises in convolutional architectures equipped with residual parameterizations. Residual convolutional networks (he2016resnet) introduce an explicit near–identity accumulation mechanism across depth, which aligns closely with the structural setting analyzed in Section 4. In this case, log–shift invariance of the renormalized shell couplings can be established under additional assumptions, rather than merely postulated. The residual parameterizations, as introduced in ResNet architectures (he2016resnet), have also been shown to promote stable Jacobian spectra across depth (tarnowski2019dynamical).
@@ 6.3 Transformer architectures
Transformer architectures combine multiple computational modules within each block and rely heavily on normalization and residual connections. These design choices promote stable training dynamics and help maintain bounded Jacobian norms over long training horizons, consistent with Condition 3 (vaswani2017attention; ba2016layernorm).
@@ 6.4 Structured state–space models
Structured state–space models and related architectures introduce gu2021efficiently; smith2022simplified; gu2024mamba explicit recurrence or long–range temporal structure. From the perspective of Condition 1, such models are compatible with the sufficient conditions provided that recurrent interactions remain effectively contractive, so that Jacobian evolution does not induce unbounded long–range coupling.

### v4

@@ 1 Introduction
Empirical power–law scaling has emerged as one of the most robust and reproducible phenomena in modern deep learning. Across model families, tasks, and training regimes, performance metrics such as loss, error, or perplexity often exhibit smooth power–law dependence on model size, dataset size, or compute budget (Hestness et al., 2017; Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022). More recently, scaling behavior has been observed to persist beyond classical regimes, extending to precision, data pruning, and architectural interventions (Sorscher et al., 2022; Kumar et al., 2024). Despite its empirical ubiquity, the theoretical origins of these power laws remain only partially understood.
A growing body of work has sought to explain neural scaling laws through kernel limits and mean–field approximations (Jacot et al., 2018; Lee et al., 2019; Yang, 2021), as well as through dynamical models that interpolate between lazy and feature–learning regimes (Bordelon et al., 2024; Bordelon and Pehlevan, 2024; Dominé et al., 2024). While these approaches provide valuable insight into specific limits or architectures, they do not by themselves explain why power–law behavior arises so broadly, nor under what conditions it should be expected to fail. In particular, empirical evidence clearly indicates that scaling laws do not hold universally: depending on architecture, optimization stability, data distribution, or training regime, scaling behavior may degrade, break, or transition between distinct regimes (Sorscher et al., 2022; Keskar et al., 2017; Ghorbani et al., 2019).
In recent work, the Generalized Resolution–Shell Dynamics (GRSD) framework was proposed as a spectral and operator–theoretic description of learning dynamics in deep networks (Zhang, 2025a). GRSD formulates learning as an energy transport process across logarithmic spectral shells, leading to a coarse–grained, one–dimensional conservation law for error energy. Within this framework, empirical neural scaling laws correspond to a particularly simple form of the renormalized shell velocity: a power–law function of the spectral coordinate. However, in GRSD this power–law form is not automatic. Rather, it reflects a nontrivial structural property of the learning process—namely, that the shell dynamics admits a renormalizable closure in the sense familiar from statistical physics and turbulence theory (Kadanoff, 1966; Wilson, 1983; Kolmogorov, 1995; Forster et al., 1977; Spohn, 2012).
@@ 2 Generalized Resolution–Shell Dynamics
In this section we briefly review the Generalized Resolution–Shell Dynamics (GRSD) framework, which provides the dynamical and spectral setting for the results of this paper. Our presentation is intentionally concise and focuses only on the aspects of GRSD that are required to formulate and analyze power–law renormalizability. We refer to (Zhang, 2025b) for a detailed and comprehensive treatment.
@@ 2.1 Spectral shells and shell energies
which plays the role of a time-dependent kernel governing error dynamics. Operators of this form have been extensively studied in learning theory and operator-based analyses of neural networks (Rosasco et al., 2010; Koltchinskii and Lounici, 2017).
@@ 2.2 Renormalized shell dynamics
where $J(\lambda,t)$ is a spectral flux density describing energy transport across scales, and $D(\lambda,t)$ collects dissipative contributions. This equation is the learning-theoretic analogue of energy cascade equations in turbulence and large-scale interacting systems (Kolmogorov, 1995; Forster et al., 1977; Spohn, 2012).
@@ 3.1 Learning configurations and Jacobian dynamics
which governs the instantaneous learning dynamics in function space. Operator evolutions of this form arise naturally in both kernel limits and feature-learning regimes (Jacot et al., 2018; Lee et al., 2019; Bordelon et al., 2024; Zhang, 2025b).
@@ 3.2 Statement of sufficient conditions
Assumptions of this type—uniformly bounded Jacobians/gradients along the optimization trajectory, sometimes supplemented by smoothness controls—are standard in analyses of overparameterized learning and trainability; see, e.g., bounded-Jacobian conditions in Terjék (2022) and Jacobian-spectrum stability via dynamical isometry in Pennington et al. (2017); Tarnowski et al. (2019).
@@ 3.3 Main theorem
Fix a learning configuration and consider the induced GRSD shell dynamics defined on logarithmic spectral shells. Suppose Conditions 14 hold on a training horizon $t\in[t_{0},T]$. Assume in addition the standard GRSD structural property that intra–shell couplings are antisymmetric, so that shell–internal transfers cancel in the energy balance.
@@
is uniquely constrained to the form
$v(\lambda,t)=\frac{c_{0}}{t}\lambda,\qquad t\in[t_{0},T],$
@@
When a learning–rate schedule $\eta(t)$ is present, the parameter dynamics take the form $\dot{\theta}=-\eta(t)\nabla_{\theta}\mathcal{L}$ and can be mapped to the above gradient–flow form by introducing a reparameterized time Li et al.
@@ 3.4 Proof overview
We briefly outline the structure of the proof; technical details are deferred to the appendix. First, Conditions 13 imply that locality and summable incoherence of Jacobian blocks propagate along the training trajectory. At the level of the error dynamics $\dot{e}=-Me$ with $M=JJ^{*}$, a Taylor expansion around a fixed log–shell basis shows that direct off-diagonal error exchange between shells separated by more than one shell index arises only at second order in variations of $M$ Tian (2025). Consequently, the leading–order shell dynamics is governed solely by nearest–neighbor boundary fluxes, while nonlocal contributions enter only as subleading renormalizations (see Appendix A.2).
Second, Condition 4 implies homogeneity of shell–averaged energy and flux densities under spectral rescaling. Finally, homogeneity of the ratio $v(\lambda,t)=J(\lambda,t)/\varepsilon(\lambda,t)$ enforces a power–law functional form. Detailed proofs are provided in Appendix A.2.
@@ 4.2 Residual learning implies log–shift invariance
The residual form $A_{k}=I+\varepsilon G_{\ell}$ induces additive log–spectral increments across depth. Under the absence of depth–wise parameter sharing and the controlled path regularity of Condition 3, these increments define a stationary Markov–additive process on spectral directions. Two mechanisms govern the depth threshold. First, Berry–Esseen type bounds Bolthausen (1982) yield self–averaging of accumulated log–spectral increments once $L\gtrsim\varepsilon^{-2}\eta^{-2}$. Second, the near–identity residual perturbations induce quantitative mixing on the projective space, leading to isotropization of spectral directions on a scale $L\gtrsim\varepsilon^{-2}\log(1/\eta)$. Beyond the larger of these two thresholds, bin–averaged coupling statistics become translation–invariant in logarithmic scale. A complete proof is provided in Appendix A.3. ∎
Theorem 2 identifies residual learning as a mechanism by which individual layerwise Jacobian contributions become asymptotically free of absolute scale information once depth exceeds a finite threshold. Its proof relies in an essential way on the near–identity residual parameterization and depth–wise statistical stationarity. The emergence of directionally homogeneous behavior under repeated small random perturbations is consistent with classical results on random matrix products and Markovian mixing on projective spaces (Furstenberg and Kesten, 1960; Levin et al., 2009). As a result, the argument does not directly extend to architectures with explicit gating or depth–dependent weighting, nor to general non–skip architectures. This limitation should not be interpreted as ruling out scaling behavior in such models; rather, it indicates that different analytical tools would be required to verify Condition 4 outside the residual setting.
@@ 5.1 Boundedness of the gradient computation graph
A canonical example of a bounded computation graph is a directed acyclic graph (DAG) Abadi et al. (2016); Paszke et al. (2019), such as those induced by standard feedforward networks, residual architectures, or transformers. In these settings, gradient propagation is naturally confined by the depth of the graph, and the influence of any parameter update cannot spread arbitrarily far in an infinitesimal time step. As a result, the induced Jacobian evolution is uniformly bounded in the sense required by Condition 1.
By contrast, recurrent architectures with unconstrained feedback loops may violate this boundedness condition. In particular, when cyclic structures contribute divergent amplification—such as repeated multiplication by unstable recurrent weights—the Jacobian evolution may accumulate unbounded long-range couplings. Classical recurrent neural networks (RNN) Hopfield (1982) and certain long short-term memory (LSTM) Hochreiter and Schmidhuber (1997) configurations fall into this category when their recurrent dynamics is not properly controlled. In such cases, the GRSD coarse-graining procedure may fail, as infinitesimal updates can induce global spectral reorganization.
Importantly, the presence of cycles alone does not preclude boundedness. Recurrent or iterative architectures may still satisfy Condition 1 provided that the contribution of cycles is contractive or effectively damped. For example, when recurrent interactions decay exponentially Gu et al. (2021); Smith et al. (2022); Gu and Dao (2024); Peng et al. (2023); Xu et al. (2025), the cumulative influence of long feedback paths remains finite, and Jacobian evolution can be bounded despite the existence of loops. From the GRSD perspective, such architectures behave similarly to bounded-depth computation graphs at the level of infinitesimal learning dynamics. We provide a precise proposition and a proof sketch for this case in Appendix 2
@@ 5.2 Initial functional incoherence
Condition 2 constrains the geometry of the Jacobian at initialization. It requires that correlations between distant Jacobian blocks, measured through operator inner products, decay sufficiently fast so that their tail is summable. Modern zero-mean random i.i.d. initializations Glorot and Bengio (2010); He et al. (2015); Klambauer et al. (2017) commonly used in deep learning naturally satisfy the condition of initial gradient incoherence, in the sense that gradient components associated with distinct functional modes are uncorrelated at initialization.
@@ 5.3 Controlled Jacobian evolution
This assumption should be viewed as an abstraction of common stability-inducing practices in modern deep learning, such as normalization, residual connections, and conservative optimization schedules (Terjék, 2022; Pennington et al., 2017; Tarnowski et al., 2019). However, the condition itself is agnostic to the specific optimization algorithm and does not rely on stochasticity or noise-induced regularization.
@@ 5.4 Log–shift invariance and rigidity of renormalized dynamics
Within the GRSD framework, log–shift invariance alone enforces homogeneity of the renormalized shell equations, but does not by itself fix the functional form of the shell velocity. The emergence of a power–law velocity is instead a rigidity phenomenon: when log–shift invariance is combined with the intrinsic time–rescaling covariance of gradient flow, all admissible functional forms except pure power laws are excluded. Such rigidity phenomena, in which scale invariance combined with consistency constraints restrict admissible large-scale behavior to power laws, are well known in renormalization-group analyses of critical phenomena and turbulent transport (Kadanoff, 1966; Wilson, 1983; Forster et al., 1977). If Condition 4 fails, the shell dynamics may exhibit scale–dependent drift, moving spectral reference points, or broken scaling behavior, even if the remaining conditions are satisfied.
@@ 6.1 Multilayer perceptrons
Multilayer perceptrons (MLPs) Rosenblatt (1957) provide the simplest setting in which the sufficient conditions can be examined. Their feedforward structure naturally induces locality in the gradient computation graph, consistent with Condition 1. Moreover, standard random initialization schemes typically lead to weak functional correlations between distant layers in sufficiently wide networks, aligning with Condition 2 in early training regimes.
@@ 6.2 Convolutional networks
An important exception arises in convolutional architectures equipped with residual parameterizations. Residual convolutional networks (He et al., 2016) introduce an explicit near–identity accumulation mechanism across depth, which aligns closely with the structural setting analyzed in Section 4. In this case, log–shift invariance of the renormalized shell couplings can be established under additional assumptions, rather than merely postulated. The residual parameterizations, as introduced in ResNet architectures (He et al., 2016), have also been shown to promote stable Jacobian spectra across depth (Tarnowski et al., 2019).
@@ 6.3 Transformer architectures
Transformer architectures combine multiple computational modules within each block and rely heavily on normalization and residual connections. These design choices promote stable training dynamics and help maintain bounded Jacobian norms over long training horizons, consistent with Condition 3 (Vaswani et al., 2017; Ba et al., 2016).
@@ 6.4 Structured state–space models
Structured state–space models and related architectures introduce Gu et al. (2021); Smith et al. (2022); Gu and Dao (2024) explicit recurrence or long–range temporal structure. From the perspective of Condition 1, such models are compatible with the sufficient conditions provided that recurrent interactions remain effectively contractive, so that Jacobian evolution does not induce unbounded long–range coupling.
@@ 6.5 Summary
7 Experimental Validation of Condition 4
The theory developed in this work relies on three intertwined empirical assumptions about the spectral organization of learning dynamics:
  1. 1.
    Controlled inter-shell coupling. Spectral shells interact weakly, so that cross-shell transport acts as a subleading and stable perturbation.
  2. 2.
    Necessity of shell coarse-graining. Intra-shell mixing is strong compared to inter-shell transport, making coarse-graining over spectral shells essential for any effective description.
  3. 3.
    Realization of Condition 4 by residual learning. In residual architectures, the induced spectral transport operator becomes approximately Toeplitz in an intermediate spectral window, reflecting approximate shift-invariance across shells.
In this section, we provide minimal yet direct empirical evidence for all three claims using controlled experiments on CIFAR-10 Krizhevsky et al. (2014). We emphasize that the goal is not to exhaustively characterize training dynamics, but to validate the structural assumptions required for a closed, shell-based description of learning.
7.1 Experimental setup
We evaluate the spectral coupling structure induced by training dynamics on image classification using CIFAR-10. All experiments are conducted with ResNet-18 and a corresponding plain (non-residual) architecture with identical depth and channel configuration.
Models.
The residual model is a standard ResNet-18. The plain counterpart (PlainNet-18) is constructed by removing all residual skip connections while keeping convolutional blocks, batch normalization, and parameter counts identical. Both models share the same initialization and trained checkpoints.
Dataset and checkpoints.
We use the CIFAR-10 validation set with batch size $B=64$. Unless otherwise stated, all spectral operators are evaluated at a fixed training checkpoint using a local finite-difference approximation of the operator derivative (see below), ensuring that both architectures are compared at the same effective training stage.
Error representation.
The error vector $e(x;\theta)$ is constructed from model logits using the $p-\text{onehot}$ formulation,
$e=p_{\theta}(x)-\mathrm{onehot}(y),$
flattened over batch and class dimensions. This choice aligns the error space with the Fisher-like geometry of classification loss and is used consistently across all experiments.
7.2 Spectral operator estimation
Our goal is to probe the spectral structure of the operator
$M(\theta)=\mathbb{E}_{x}\big[J(x;\theta)J(x;\theta)^{\top}\big],$
where $J=\partial e/\partial\theta$ is the Jacobian of the error vector with respect to model parameters.
Monte Carlo operator evaluation.
The expectation defining $M$ is approximated using a small number ($S=4$) of fixed mini-batches sampled from the validation set. Matrix–vector products $Mv$ are computed without explicitly forming $M$, using a Jacobian–vector product followed by a vector–Jacobian product (JVP–VJP), enabling scalable operator access.
Block Lanczos projection.
We apply a block Lanczos algorithm Golub and Meurant (2010); Ubaru et al. (2017) with block size $b=8$ and $k=16$ steps, resulting in a Krylov subspace of dimension $r=bk=128$. This produces a block-tridiagonal matrix $T$ whose eigenpairs $\{(\lambda_{i},u_{i})\}$ approximate the spectrum of $M$ in the probed subspace.
Directional derivative of the operator.
To probe spectral transport, we compute a finite-difference approximation of the operator derivative
$\dot{M}\approx\frac{M(\theta-\varepsilon\nabla\mathcal{L})-M(\theta)}{\varepsilon},$
where $\varepsilon$ is a small step along the gradient direction. The projected operator $\Omega=U^{\top}(Q^{\top}\dot{M}Q)U$ is expressed in the Ritz eigenbasis of $M$.
More details regarding the estimation algorithm are included in Appendix B.
7.3 Shell binning and Toeplitz diagnostics
Eigenvalues of $M$ are binned into logarithmically spaced spectral shells based on $\log\lambda$. Given the projected operator $\Omega$, we define a coarse-grained shell coupling matrix $K$ by averaging squared couplings within each shell pair:
$K_{ij}=\mathbb{E}_{u\in\mathcal{S}_{i},v\in\mathcal{S}_{j}}\big[|\Omega_{uv}|^{2}\big].$
Toeplitz residual.
To quantify the degree of shift-invariance across shells, we compare $K$ to its best Toeplitz approximation obtained by averaging along diagonals. The normalized Frobenius residual
$\mathrm{Res}(K)=\frac{\|K-\mathcal{T}(K)\|_{F}^{2}}{\|K\|_{F}^{2}}$
serves as our primary diagnostic: lower values indicate stronger Toeplitz structure i.e. the coupling strength relies on the relative scale shift rather than absolute scales.
All scores are computed over sliding spectral windows. We focus on windows centered in the intermediate spectral regime, where the assumptions underlying our effective theory are expected to hold.
7.4 Residual learning induces Toeplitz spectral structure
Figure 1 compares the Toeplitz residual of ResNet-18 and PlainNet-18 as the analysis window moves from the spectral edge toward the spectral center.
Residual networks exhibit a rapid and monotonic improvement in Toeplitz structure when approaching the intermediate spectral regime. In contrast, the plain architecture shows significantly weaker Toeplitz behavior with little sensitivity to window position. This provides direct empirical evidence that residual learning facilitates the approximate shift-invariance across spectral shells required by Condition 4.
Near the spectral edges, Toeplitz structure deteriorates for both models. This behavior is expected, as these regions lie outside the effective spectral and time window assumed by the theory, rather than indicating a failure of the condition itself.
Figure 1: Toeplitz residual of the coarse-grained spectral coupling matrix as a function of distance to the spectral edge (larger values correspond to windows closer to the intermediate spectral regime). Residual networks exhibit a rapid and monotonic improvement in Toeplitz-ness toward the spectral center, while the corresponding plain networks show substantially weaker and flatter behavior.
7.5 Controlled inter-shell coupling and shell coarse-graining
To further dissect the observed Toeplitz structure, we separate contributions from intra-shell (on-diagonal) and inter-shell (off-diagonal) couplings.
Figure 2 shows that inter-shell couplings are suppressed by several orders of magnitude toward the spectral center, while intra-shell mixing remains comparatively large and stable. This clear separation indicates that cross-shell interactions act as a controlled perturbation, consistent with the presence of effective spectral gaps between shells.
These results justify the shell-based coarse-graining used throughout our theoretical analysis: rapid intra-shell equilibration dominates local dynamics, while inter-shell transport governs slower, structured spectral flow.
Figure 2: Comparison between off-diagonal (inter-shell) and on-diagonal (intra-shell) Toeplitz residuals as a function of spectral window position. Inter-shell couplings are strongly suppressed toward the spectral center, while intra-shell mixing remains comparatively large and stable.